# Explaining mountain pine beetle dynamics: From life history traits to large, episodic outbreaks

### v1

@@ 2.5 Causal analysis
When the simulation model is instantiated with realistic parameter values, the five focal life-history traits are turned on. To turn the life-history traits off, we alter specific parameter values.
To turn off the Allee effect, we set the Allee threshold, $a$, to infinity. As a consequence, the probability of a successful attack is always one, or equivalently, the number of initial attacks equals the number of successful attacks. To turn off stand-dependent dispersal, we set $\lambda_{0}$ to infinity, such that the $s_{y}=s_{0}$ and $\lambda_{y}=1$ in Eq. 2 & Eq. 3 respectively. To turn off the size-dependent attack preference, the piecewise linear function for attack preference, Eq. 7, becomes a step function,
@@
To turn off size-dependent fecundity, we set $\zeta_{1}=0$ and set $\zeta_{0}$ to the average emergence density (across killed tree surface area in the Klein et al., dataset), which is calculated as
@@
Finally, we turn off time-scale separation by forcing lodgepole pine to grow quickly. To accomplish this, we iterate pine mortality, growth, and recruitment (Eq. 12–Eq. 14) 50 times between bouts of beetle attacks, such that year-old trees have 8-inch diameters.
We create simulation scenarios where either 0, 1, or 2 of the life history traits have been turned off. The result is 10 simulation scenarios: 9 counterfactual and 1 realistic (i.e., all of the traits are turned on). For each scenario, we draw 1,000 sets of parameters from the joint posterior distribution of the statistical models. For each parameter set, we simulate 3,000 years of data, starting with zero beetles and equilibrium forest conditions.
@@ 2.6 Describing population dynamics
    Peak beetle density. For each simulated time series, we take the third quartile of all peak values of $B_{y}$, the density of emerging female beetles. The third quartile represents a typical large outbreak, which is relevant because large outbreaks are likely the driver of range expansion (see the Discussion, Section 4).
@@
    An observation-based estimate is calculated using British Columbia’s Aerial Overview Survey data from 1990-2021. First, we estimate the number of infested trees using the methodology in Fig. D.3 caption, aggregate data into $1\text{km}^{2}$ pixels, and subset to locations with moderate-to-severe outbreaks (cumulative infestations in the upper quartile). Next, we calculate the difference between the year with the maximum number of infested trees and the year before infested trees were first detected. Finally, we take the mean of this quantity across all locations.
@@
The five metrics above can be used heuristically to classify simulated time series into qualitative classes of population dynamics: episodic boom-bust dynamics, population cycles, or equilibrium dynamics. If the proportion of time spent in the endemic states is $\geq 1/2$, then the time series has Boom-bust dynamics. If the proportion of time spent in the endemic states is $<1/2$, and the outbreak period is $<50$ years, then the time series has Cycles. If the proportion of time spent in the endemic states is $<1/2$, and the outbreak period is $\geq 50$ years, then the time series has Equilibrium dynamics.
@@ 3 Results
The statistical model successfully predicts in-sample data (Fig. 14, panel A) and out-of-sample data. The exception is that the out-of-sample predictive intervals for 1970 are wide (Fig. 14, panel B), which upon further investigation, was found to be the result of uncertain estimates of the stand-dependent dispersal parameters, $\lambda_{0}$ and $\lambda_{1}$ (Fig. 14, panel C). The estimation of these parameters is highly dependent on the data from 1969–1970. This would be indicative of overfitting (i.e., fitting the noise) if the decrease in attack density in 1970 was the result of noise, rather than stand-dependent dispersal. The obvious source of statistical noise is the weather, but the weather leading up to the summer of 1970 was unexceptional (Figs. D.5 & D.4). Taken together, evidence for the presence of stand-dependent dispersal (Section 1.2.2) and the absence of anomalous weather patterns imply that the model did not overfit.
Various diagnostics showed that the statistical models converged to the global posterior distribution, and that nearly all prior distributions had a negligible impact on the results (Appendix B.1). The only exception was the Allee threshold parameter, whose marginal posterior variance was similar in magnitude to its marginal prior variance, indicating that the Klein et al., data is not informative with respect to the Allee effect. This is unsurprising, as discerning the Allee effect is challenging without data from trees that were attacked but not successfully killed (see Section Section 1.2.1), data which is absent in Klein et al.,’s 1978 publication. However, the Allee threshold estimate is not devoid of information, since the prior distribution was derived from the Allee effect sub-model fit to the Boone et al., dataset (see Table C.2 & Appendix A.6).
Figure 14: Observations of successful attack density in the Klein et al., dataset (operationalized as gallery starts), with superimposed model-based predictive intervals (PIs). Panel A shows in-sample predictions for the full statistical model; panel B shows out-of-sample, leave-one-year-out (LOYO) predictions, where the model is not trained on the focal year’s data; and panel C shows LOYO predictions when the stand-dependent dispersal parameters ($\lambda_{0}$ and $\lambda_{1}$) are fixed at the medians of their respective marginal posterior distributions (see Table 1). Panels B & C demonstrate that data from a single beetle generation (1969–1970) has an overwhelming influence on the estimation of $\lambda_{0}$ and $\lambda_{1}$, but that all other parameter estimates are robust.
@@
The statistical models also produced reasonable parameter estimates. In particular, the posterior distributions of attack density and emergence density (from a 10-inch diameter tree) are consistent with previous estimates (Figures 15 & 16). The mean overstory density, 338 trees per acre or 835 trees per hectare, agrees with estimates from regenerating stands in southern British Columbia (Nigh et al.,, 2008, mean $=$ 893 stems/ha) and lodgepole-dominated stands in the U.S. Rocky Mountains (Pfeifer et al.,, 2011; Meddens et al.,, 2011, 2012; mean $=$ 1065 stems/ha). The minimum susceptible tree diameter is estimated at 7.34 inches, which is consistent with the often-cited 8-inch (or 20 cm) threshold (Hopping and Beall,, 1948; Cole,, 1969; Safranyik,, 1988). The baseline dispersal mortality (i.e., 1-$s_{0}$) is 0.55, which is consistent with the 0.5–0.7 range provided by previous studies of bark beetles (Schmid,, 1977; Safranyik,, 1989; Latty,, 2007; Pope et al.,, 1980).
Figure 15: Frequency distributions of attack density. The Model posterior category gives the marginal posterior of $1/\gamma$ from our statistical model of the Klein et al., outbreak. The AB disk data category contains the within-site means of attack density from 300 different sites across Alberta (author N/A,, 2020). The Other studies category includes estimates obtained via entry holes or gallery starts from a number of studies, mostly United States Forest Service report: Tishmack et al., (2005); Mccambridge, (1967); Negrón, (2018); Reid, (1963); Raffa and Berryman, (1983); Rasmussen, (1980); Shepard, (1965); Knight, (1959); De Leon, (1939); Safranyik, (1988); Safranyik and Vithayasai, (1971); Safranyik, (1968); Parker, (1979); Schmid, (1972); Blackman, (1931); Whiteside, (1937); Washburn and Cole, (1959); author N/A, (1963); Peterman, (1974); Klein et al., (1978); Safranyik and Linton, (1991).
Figure 16: Frequency distributions of female emergence density. Females are estimated with (2/3)×(3/2)×(2/3$\times$(3/2$\times$( 2 / 3 ) × ( 3 / 2 ) × emergence holes, or with (2/3)×(2/3$\times$( 2 / 3 ) × emerging beetles or spring brood size. Note that (2/3) is the typical proportion of females in a brood (Reid,, 1962; Cole et al.,, 1976), and (3/2) is the average number of beetles emerging from each hole (Safranyik and Linton,, 1985; Peterman,, 1974, Fig. 13). The Model posterior category gives the marginal posterior of $\zeta_{0}+\zeta_{1}*10$, the emergence density on a 10-inch DBH tree; 10 inches is approximately the average DBH of MPB-killed trees in the Klein et al., dataset. The AB disk data category contains estimates of emergence from 300 different sites across Alberta (author N/A,, 2020). The Other studies category includes estimates from a number of studies, mostly United States Forest Service reports: Tishmack et al., (2005); Negrón, (2018); Reid, (1963); Raffa and Berryman, (1983); Rasmussen, (1980); De Leon, (1939); Safranyik, (1988); Cole, (1969); Schmid, (1972); Whiteside, (1937); Beal, (1938); Peterman, (1974); Klein et al., (1978); Safranyik and Linton, (1985, 1991).
@@
a Allee threshold 392 331 12.1 1211.68175
$s_{0}$ max dispersal survival 0.474 0.0676 0.363 0.627
$\lambda_{0}$ dispersal threshold 0.652 0.0339 0.582 0.705
@@
$-\beta_{0}/\beta_{1}$ susc. DBH threshold 7.43 0.643 6.04 8.43
$1/\gamma$ attack density 66.2 3.78 59.2 74.1
$\zeta_{0}+\zeta_{1}*10$ emerge density (10 in DBH) 244 25.4 192 297
$K-\sum_{j=1}^{18}(1-m)^{(j-1)}mK$ Overstory density 338 86.0 193 529
@@
Table 1: Summary statistics of marginal posterior distributions for model parameters, including the Mean, standard deviation, and credible intervals ($\text{CI}_{2.5\%}$, $\text{CI}_{97.5\%}$). The table also includes derived quantities, including the diameter at breast height (DBH) of the smallest susceptible trees; the model-based estimate of successful attack density (gallery starts per square meter); the emergence density (females per square meter) for 10-inch DBH tree; and the density of overstory trees (stems per acre). Overstory trees are operationalized with a standard cutoff: DBH $\geq 7\text{cm}\approx 2.76\text{in}$. This explains the summation from 0 to 18 in the formula: 2.76in cutoff0.157in per $\frac{2.76\text{in cutoff}}{0.157\text{in per year}}\approx 18\;\text{years}$. Finally, the table contains fixed parameters, whose values were obtained via literature search. For more elaborate descriptions and parameter units, see supplementary Table A.1. For the summary statistics of sub-model noise parameters, see supplementary Table C.1.
@@
Given our understanding of MPB dynamics, three life-history traits are necessary for episodic outbreaks: Stand-dependent dispersal, an Allee effect, and time-scale separation (Fig. 19; Table 2, column Boom-Bust). Without the stand-dependent dispersal, MPB and lodgepole pine attain a noisy equilibrium where there is a consistently low density of beetles (Fig. 19C), but not low enough to technically be endemic. Without the Allee effect, a classic population cycle appears (Fig. 19D). Without time-scale separation, MPB and lodgepole pine attain a noisy equilibrium; in contrast with the no-stand-dependent-dispersal scenario, this equilibrium sustains large numbers of beetles and susceptible trees (Fig. 19E).
The inclusion of stand-dependent dispersal leads to better predictions of outbreak magnitude and large-tree density (Fig. 17). The Allee effect leads to better predictions of the outbreak recurrence time and growth phase length. Specifically, without the Allee effect, the cycles occur too frequently and the growth phase of each cycle is too long. Stand-dependent dispersal primarily explains the fact that large trees remain despite MPB outbreaks.
@@
The two remaining life-history traits — size-dependent tree preference and size-dependent fecundity — have no effect on the qualitative nature of population dynamics, i.e., whether a simulation was classified as Boom-Bust or Cycles. However, they do have quantitative effects on some of the fine-grained properties of outbreaks: the combination of both size-dependent tree preference and size-dependent fecundity lead to a 25% increase in the peak beetle density (from 16 to 20 thousand) and a 10% decrease in the length of growth phase (from 5.8 years to 5.2 years). In other words, the combination of these life-history traits makes outbreaks moderately more explosive and slightly more rapid.
@@
    Life history traits     Has stand-dependent dispersal     Has Allee effect     Has size-dependent preference     Has size-dependent fecundity     Has time-scale separation     Metrics Mean $\pm$ (SD)     Peak beetle density, 1000s of emerging females per acre     Outbreak growth phase, years     Outbreak period, years     Large tree density, trees $\geq$ 8 in DBH per acre     Prop. time endemic     Prop. sims without peaks     Prop. sims without outbreaks     Probability of Dynamics     Boom-Bust     Cycles     Equilibrium
20 (8.7) 5.2 (1.1) 44 (130) 44 (9.6) 0.66 (0.16) 0.00 (0.00) 0.00 0.86 0.10 0.03
9.5 (10) 7.2 (1.1) 179 (251) 10 (12) 0.16 (0.24) 0.01 (0.09) 0.44 0.16 0.05 0.79
6.6 (3.2) 9.4 (2.3) 28 (63) 38 (9.9) 0.07 (0.15) 0.00 (0.00) 0.04 0.05 0.83 0.13
16 (7.2) 5.8 (1.4) 51 (57) 43 (9.3) 0.66 (0.16) 0.00 (0.00) 0.00 0.87 0.09 0.03
16 (6.7) 5.8 (1.3) 60 (184) 45 (10) 0.65 (0.16) 0.01 (0.08) 0.01 0.86 0.10 0.03
73 (24) NA 4.8 (0.57) 46 (11) 0.00 (0.03) 0.00 (0.00) 0.99 0.00 1.0 0.00
3.3 (0.88) 15 (1.8) 278 (378) 2.4 (4.2) 0.00 (0.00) 0.01 (0.09) 1.0 0.00 0.06 0.94
11 (13) 8.5 (1.3) 151 (168) 11 (14) 0.18 (0.26) 0.00 (0.06) 0.43 0.18 0.07 0.75
4.5 (2.2) 9.2 (2.2) 1078 (838) 3.8 (8.9) 0.03 (0.13) 0.08 (0.27) 0.87 0.03 0.00 0.97
71 (17) NA 4.7 (0.12) 44 (8.6) 0.00 (0.00) 0.00 (0.00) 1.0 0.00 1.0 0.00
5.6 (1.9) 10 (2.6) 34 (124) 37 (9.5) 0.06 (0.12) 0.00 (0.00) 0.04 0.02 0.78 0.21
5.9 (2.6) 10 (2.7) 34 (127) 39 (11) 0.05 (0.13) 0.00 (0.00) 0.16 0.03 0.78 0.19
@@
15 (6.8) 5.8 (1.3) 57 (110) 43 (9.8) 0.66 (0.17) 0.00 (0.00) 0.00 0.87 0.08 0.05
71 (13) NA 4.2 (0.33) 41 (6.8) 0.01 (0.09) 0.01 (0.10) 1.0 0.01 0.99 0.00
68 (16) NA 4.9 (0.19) 46 (9.2) 0.01 (0.09) 0.01 (0.10) 1.0 0.01 0.99 0.00
Table 2: Results of the simulation study. A check mark indicates that a life history trait is present in the simulation scenario. The quantitative values take the format of “Mean $\pm$ (SD)”, where both summary statistics are calculated across the posterior distribution. In addition to the 5 metric of population dynamics (presented in Section 2.6), there are two extra metrics: 1) the proportion of simulations where neither peaks nor troughs were detected after applying the moving-average smoothing and various thresholds described in Section 2.6; 2) the proportion of simulations where no outbreaks were detected, where an outbreak is defined as a trough $\rightarrow$ peak $\rightarrow$ trough sequence where the population attains the endemic state ($<16$ females per acre) before and after the peak. If a simulation scenario produced both NA and non-NA values for a particular metric, then the NA values were omitted when calculating that metric.
Figure 17: A comparison of population dynamic metrics for various combinations of stand-dependent dispersal and the Allee effect. The histograms show the frequency of predicted values across posterior samples. The vertical lines show observation-based estimates from the Klein et al., dataset or independent sources. Simulation scenarios without stand-dependent dispersal resulted in few peaks and outbreaks (see Table 2), hence the “N/A” label. The five metrics and the observation-based estimates are explained in Section 2.6.
Figure 18: The posterior probability of various types of population dynamics, across various simulation scenarios where stand-dependent dispersal and the Allee effect have been turned on or off.
Figure 19: A time series of Aerial Overview Survey data, along with representative time series from various simulation scenarios.

### v2

@@ 2.4 Causal analysis
When the simulation model is instantiated with realistic parameter values, the five focal biological features are turned on. To turn the features off, we alter specific parameter values.
To turn off the Allee effect, we set the Allee threshold, $a$, to infinity. As a consequence, the probability of a successful attack is always one, or equivalently, the number of initial attacks equals the number of successful attacks. To turn off stand-dependent dispersal mortality, we set $\lambda_{0}$ to infinity, such that the $s_{y}=s_{0}$ in Eq. 35. To turn off the large-tree preference, the piecewise linear function for attack preference, Eq. 4, becomes a step function,
@@
To turn off tree size-dependent fecundity, we set $\zeta_{1}=0$ and set $\zeta_{0}$ to the average emergence density (across killed tree surface area in the Klein et al., dataset), which is calculated as
@@
Finally, we turn off time-scale separation by forcing lodgepole pine to grow quickly. To accomplish this, we iterate density-independent pine mortality, growth, and recruitment (Eq. 10–Eq. 11) 50 times between bouts of beetle attacks, such that year-old trees have 8-inch diameters.
We create simulation scenarios where either 0, 1, or 2 of features have been turned off. The result is 10 simulation scenarios: 9 counterfactual and 1 realistic (i.e., all of the features are turned on). For each scenario, we draw 1,000 sets of parameters from the joint posterior distribution of the statistical models. For each parameter set, we simulate 3,000 years of data, starting with zero beetles and equilibrium forest conditions.
@@ 2.5 Describing population dynamics
    Peak beetle density. For each simulated time series, we take the third quartile of all peak values of $B_{y}$, the density of emerging female beetles. The third quartile represents a typical large outbreak, which is relevant because large local outbreaks disproportionaly contribute to spatial spread and range expansion (see the Discussion, Section 4).
@@
    An observation-based estimate is calculated using British Columbia’s Aerial Overview Survey data from 1990-2021. First, we estimate the number of infested trees using the methodology in Fig. E.1 caption, aggregate data into $1\text{km}^{2}$ pixels, and subset to locations with moderate-to-severe outbreaks (cumulative infestations in the upper quartile). Next, we calculate the difference between the year with the maximum number of infested trees and the year before infested trees were first detected. Finally, we take the mean of this quantity across all locations.
@@
The five metrics above can be used heuristically to classify simulated time series wtih respect to their population dynamical behavior: excitable dynamics (i.e., episodic boom-bust dynamics), population cycles, or equilibrium dynamics. If the proportion of time spent in the endemic states is $\geq 1/2$, then the time series has excitable dynamics. If the proportion of time spent in the endemic states is $<1/2$, and the outbreak period is $<50$ years, then the time series has Cycles. If the proportion of time spent in the endemic states is $<1/2$, and the outbreak period is $\geq 50$ years, then the time series has Equilibrium dynamics.
@@ 3 Results
3.1 Model diagnostics and parameter estimates
The statistical models performed well across multiple assessment criteria. Various diagnostics indicated that the models converged to the global posterior distribution, with most prior distributions having negligible impact on the results (Appendix A.3). The models successfully recreated the original outbreak trajectory and out-of-sample data (Appendix B.4).
@@ 3.1 Model diagnostics and parameter estimates
The parameter estimates produced by the statistical models align with previous research (Table 1). The posterior distributions of attack density and emergence density (from a 10-inch diameter tree) are consistent with over over a dozen studies across British Columbia, Alberta, and the western United States (Figures E.5 & E.6). The mean overstory density, 338 trees per acre or 835 trees per hectare, agrees with estimates from regenerating stands in southern British Columbia (Nigh et al.,, 2008, mean $=$ 893 stems/ha) and lodgepole-dominated stands in the U.S. Rocky Mountains (Pfeifer et al.,, 2011; Meddens et al.,, 2011, 2012; mean $=$ 1065 stems/ha). The minimum susceptible tree diameter is estimated at 7.46 inches, which is consistent with the often-cited 8-inch (or 20 cm) threshold (Hopping and Beall,, 1948; Cole,, 1969; Safranyik,, 1988). The baseline dispersal mortality (i.e., 1-$s_{0}$) is 0.52, which is consistent with the 0.5–0.7 range provided by previous studies of bark beetles (Schmid,, 1977; Safranyik,, 1989; Latty,, 2007; Pope et al.,, 1980). Finally, the mean attack density of 71 females per square meter in the Klein et al., dataset is close to the optimal value of 60 females per square meter, proposed by Raffa and Berryman, (1983).
@@
a Allee threshold 432 363 13.5 1338.88825
$s_{0}$ max dispersal survival 0.477 0.0701 0.363 0.639
$\lambda_{\text{resid}}$ dispersal threshold 0.652 0.0353 0.579 0.708
@@
$-\beta_{1}/\beta_{0}$ susc. DBH threshold 7.46 0.666 5.98 8.47
$1/\gamma$ attack density 66.3 3.88 59.4 74.4
$\zeta_{0}+\zeta_{1}*10$ emerge density (10 in. DBH) 244 25.1 191 296
$K-\sum_{t=0}^{17}(1-m)^{t}mK$ Overstory density 338 86.0 193 529
@@
Table 1: Summary statistics of marginal posterior distributions for model parameters, including the Mean, standard deviation, and credible intervals ($\text{CI}_{2.5\%}$, $\text{CI}_{97.5\%}$). The table also includes derived quantities, including the diameter at breast height (DBH) of the smallest susceptible trees; the model-based estimate of successful attack density (gallery starts per square meter); the emergence density (females per square meter) for 10-inch DBH tree; and the density of overstory trees (stems per acre). Overstory trees are operationalized with a standard cutoff: DBH $\geq 7\text{cm}\approx 2.76\text{in}$. This explains the summation from 0 to 18 in the formula: 2.76in cutoff0.157in per $\frac{2.76\text{in cutoff}}{0.157\text{in per year}}\approx 18\;\text{years}$. Finally, the table contains fixed parameters, whose values were obtained via literature search. For more elaborate descriptions and parameter units, see supplementary Table LABEL:tab:model_notation. For the summary statistics of sub-model noise parameters, see supplementary Table C.1.
3.2 Effects of life history on population dynamics
@@
Three key features are essential for excitable population dynamics in MPB: stand-dependent dispersal mortality, an Allee effect, and time-scale separation (Fig.,6, Fig. 7). Without the stand-dependent dispersal mortality, MPB and lodgepole pine attain a noisy equilibrium where there is a consistently low density of beetles (Fig. 6C), but not low enough to technically be endemic. Without the Allee effect, a classic population cycle appears (Fig. 6D). Without time-scale separation, MPB and lodgepole pine attain a noisy equilibrium with large numbers of beetles and susceptible trees (Fig. 6 E).
Figure 6: Time series of mountain pine beetle infestations, real vs. simulated. The real data comes from Aerial Overview Survey data in British Columbia, aggregated across a representative 1 $\text{km}^{2}$ patch. Simulated time series are representative of various simulation scenarios.
Figure 7: Stand-dependent dispersal mortality and the Allee effect lead to excitable boom-bust dynamics. Barplot proportions represent the posterior probabilities of different dynamical behaviors, classified using the methodology in Section 2.5.
The process of stand-dependent dispersal mortality explains why beetles don’t overexploit trees. It is perhaps trivial to say that if beetles disperse away and die before they have killed all the trees, then the beetles won’t kill all the trees. However, the lack of over-exploitation, or equivalently, the seemingly premature “bust” phase of the “boom-bust cycle”, is a ubiquitous feature of outbreaks (Fig. 4) and is thus worth connecting to MPB life-history. This also suggests an optimistic perspective: outbreaks are much less destructive than they theoretically could be. Stand-dependent dispersal probably occurs because chemo-sensory constraints prevent beetles from aggregating on small trees (Section 1.2.2), even though beetles have positive fitness when they do successfully attack small trees. Without stand-dependent dispersal mortality, the system would reach a stochastic equilibrium, creating a forest with few large pine trees.
Classic predator-prey theory predicts that long-lived prey should make predator-prey cycles unlikely (Murdoch et al.,, 2003). One might naively expect that the time-scale separation between beetles and trees would allow MPB to grow rapidly enough to prevent an unchecked proliferation in pine abundance, thus preventing the prey-escape phase of a population cycle (De Roos et al.,, 1990) wherein the prey species can grow unimpeded by the predator. Even though pines live much longer than beetles, we still see population cycles in the counterfactual simulations (when the Allee effect is turned off). This can be explained by the phenomenon of cohort resonance (Bjørnstad et al.,, 2004). During an outbreak, many pines are killed at roughly the same time. Subsequently, pines are recruited at the same time, and then become susceptible to MPB at the same time. Thus, there are a brief few years where the pines that MPB can “see” are growing faster than exponentially.
@@
The two remaining features — size-dependent tree preference and tree size-dependent fecundity — do not affect the qualitative nature of MPB population dynamics, i.e., whether a simulation was classified as Excitable or Cycles. However, they do have quantitative effects on some of the fine-grained properties of outbreaks: the combination of both size-dependent tree preference and tree size-dependent fecundity led to a 40% increase in the peak beetle density (from 15 to 21 thousand) and an 8% decrease in the length of growth phase (from 5.8 years to 5.4 years). In other words, the combination of these life-history traits makes outbreaks moderately more explosive and slightly more rapid.
The inclusion of stand-dependent dispersal mortality leads to better predictions of outbreak magnitude and large-tree density (Fig. 8). The Allee effect leads to better predictions of the outbreak recurrence time and growth phase length. Specifically, without the Allee effect, the cycles occur too frequently and the growth phase of each cycle is too long. Stand-dependent dispersal mortality primarily explains the fact that a portion of large trees remain despite MPB outbreaks.
@@
    Biological features     Has stand-dependent dispersal mortality     Has Allee effect     Has large-tree preference     Has tree size-dependent fecundity     Has time-scale separation     Metrics Mean $\pm$ (SD)     Peak beetle density, 1000s of emerging females per acre     Outbreak growth phase, years     Outbreak period, years     Large tree density, trees $\geq$ 8 in DBH per acre     Prop. time endemic     Prop. sims without peaks     Prop. sims without outbreaks     Probability of Dynamics     Excitable     Cycles     Equilibrium
21 (9.0) 5.4 (1.2) 44 (34) 44 (9.7) 0.68 (0.16) 0.00 (0.00) 0.00 0.90 0.08 0.01
10 (10) 7.3 (1.1) 174 (295) 11 (13) 0.18 (0.25) 0.00 (0.06) 0.44 0.18 0.05 0.78
6.4 (3.1) 9.7 (2.2) 26 (10) 38 (9.2) 0.07 (0.15) 0.00 (0.00) 0.06 0.05 0.80 0.15
16 (7.1) 5.9 (1.4) 57 (136) 43 (10) 0.66 (0.16) 0.00 (0.00) 0.00 0.89 0.07 0.04
16 (7.0) 5.9 (1.4) 60 (157) 45 (10) 0.65 (0.16) 0.00 (0.00) 0.00 0.88 0.10 0.03
72 (18) 8.0 (2.4) 4.8 (1.2) 46 (9.1) 0.00 (0.04) 0.00 (0.00) 0.99 0.00 1.0 0.00
3.3 (0.90) 13 (1.4) 293 (374) 2.6 (4.0) 0.00 (0.01) 0.01 (0.08) 1.0 0.00 0.05 0.95
11 (12) 8.5 (1.2) 221 (402) 12 (15) 0.19 (0.26) 0.01 (0.08) 0.45 0.19 0.07 0.74
4.6 (2.2) 8.9 (1.9) 1076 (876) 4.8 (9.9) 0.04 (0.15) 0.09 (0.29) 0.85 0.04 0.00 0.96
73 (16) NA 4.7 (0.12) 46 (8.8) 0.00 (0.00) 0.00 (0.00) 1.0 0.00 1.0 0.00
5.7 (2.7) 10 (2.4) 30 (17) 37 (10) 0.06 (0.14) 0.00 (0.00) 0.05 0.04 0.74 0.23
6.1 (2.9) 9.9 (2.7) 29 (36) 39 (10) 0.06 (0.15) 0.00 (0.00) 0.13 0.05 0.75 0.21
@@
15 (7.0) 5.8 (1.3) 63 (200) 42 (10.0) 0.65 (0.16) 0.00 (0.06) 0.00 0.87 0.09 0.05
76 (20) 7.4 (1.8) 9.9 (125) 42 (9.0) 0.01 (0.07) 0.00 (0.04) 0.99 0.01 0.99 0.00
66 (15) 13 (3.9) 4.9 (0.30) 46 (9.0) 0.01 (0.05) 0.00 (0.04) 0.98 0.00 1.0 0.00
Table 2: Results of the simulation study. A check mark indicates that a biological feature is present in the simulation scenario. The quantitative values take the format of “Mean $\pm$ (SD)”, where both summary statistics are calculated across the posterior distribution. In addition to the 5 metrics of population dynamics (presented in Section 2.5), there are two extra metrics: 1) the proportion of simulations where neither peaks nor troughs were detected after applying the moving-average smoothing and various thresholds described in Section 2.5; 2) the proportion of simulations where no outbreaks were detected, where an outbreak is defined as a trough $\rightarrow$ peak $\rightarrow$ trough sequence where the population attains the endemic state ($<16$ females per acre) before and after the peak. If a simulation scenario produced both NA and non-NA values for a particular metric, then the NA values were omitted when calculating that metric.
Figure 8: A comparison of population dynamic metrics for various combinations of stand-dependent dispersal mortality and the Allee effect. The histograms show the frequency of predicted values across posterior samples. The vertical lines show observation-based estimates from the Klein et al., dataset or independent sources. Simulation scenarios without stand-dependent dispersal mortality resulted in few peaks and outbreaks (see Table 2), hence the “N/A” label. The five metrics and the observation-based estimates are explained in Section 2.5.